A 20 volts 5 watts lamp is used on AC mains of 200 volts 50/pi root 11 c.p.s. Calculate the 1) capacitor, 2) inductor, to be put in series to run on its rated voltage, 3) How much pure resistance should be included in place of the above device so that the lamp can run on its rated voltage, 4) which is more economical, the capacitor, the inductor or the resistor.
Plzz..answer this question with suitable solution..

Grade:12th pass

1 Answers

Arun
4 years ago
Dear student
 
current required = P/V = 5/20 = 0.25 amp
resistance of lamp = V/I = 20/0.25 = 80 ohm
 
i)-
Z = sqrt (R^2 + 1/Wc^2)